# supplement

### v2

@@ Abstract
We compare two approaches for solving high-dimensional eigenvalue problems with low-rank structure: the inexact Lanczos method and inexact polynomial-filtered subspace iteration. Inexactness stems from low-rank compression, enabling efficient representation of high-dimensional vectors in a low-rank tensor format. A primary challenge in these methods is that standard operations, such as matrix-vector products and linear combinations, increase tensor rank, necessitating rank truncation and hence approximation. The Lanczos method constructs an approximate orthonormal Krylov basis, which is often difficult to represent accurately using low-rank tensor formats, even when the eigenvectors themselves exhibit low-rank structure. In contrast, the low-rank polynomial-filtered subspace iteration uses approximate eigenvectors (Ritz vectors) directly as a subspace basis, bypassing the need for an orthonormal Krylov basis. Our analysis and numerical experiments demonstrate that inexact subspace iteration is much more robust to rank-truncation errors compared to the inexact Lanczos method. We further demonstrate that rank-truncated subspace iteration can converge for problems where the density matrix renormalization group method (DMRG) stagnates.
keywords:
subspace iteration, Krylov methods, low-rank approximation, tensor networks
{MSCcodes}
65F15, 15A69
@@ 1 Introduction
of a Hermitian matrix $A\in\mathbb{C}^{(n_{1}\cdots n_{d})\times(n_{1}\cdots n_{d})}$ that can be written as a sum of Kronecker products
(2) $A=\sum_{j=1}^{r_{A}}A_{1}^{(j)}\otimes A_{2}^{(j)}\otimes\cdots\otimes A_{d}^{(j)},$
@@
where $r_{A}$ is the separation (or canonical polyadic) rank of $A$ and $A_{i}^{(j)}\in\mathbb{C}^{n_{i}\times n_{i}}$ for all $i=1,2,\ldots,d$ and $j=1,\ldots,r_{A}$. Examples of (2) appear naturally in quantum many-body physics and arise in many other applications from the discretization of partial differential operators on separable domains. The number of entries in $A$ and its eigenvectors $\psi\in\mathbb{C}^{n_{1}\cdots n_{d}}$ grows exponentially fast with the parameter $d$, making classical numerical methods too expensive when directly applied to $A$, even for moderate $d$. Given the tensor product structure of the operator (2), it is natural to seek eigenvectors in a compact form commensurate with the tensor product structure of the operator $A$. Many tensor networks (or formats) are available for representing operators and vectors [9], offering significant reductions in the degrees of freedom required, provided the objects can be approximated with a network of relatively small tensor rank. For quantum many-body Hamiltonians with separated eigenvalues and local interactions, it is well-known that the eigenvectors corresponding to the smallest eigenvalues can be represented exactly as low-rank tensor trains (or matrix product states) [12]. For other applications, e.g., the spatial discretization of high-dimensional elliptic operators, it has been demonstrated that eigenvectors can be well-approximated in low-rank hierarchical tensor formats [3, 10].
@@
A widely used approach for computing the extreme eigenpairs of a Hermitian matrix $A$, expressed in the form (2), is the density matrix renormalization group (DMRG) method [30, 31, 36], originally developed in the physics community. The DMRG method represents the approximate eigenvector as a tensor network and optimizes one (or two) low-dimensional tensors in the network at a time by projecting $A$ onto subspaces defined by the remaining tensors. This optimization process cycles through all tensors in the network, often multiple times, until convergence. The method achieves computational efficiency by reducing the high-dimensional eigenvalue problem to a sequence of low-dimensional eigenvalue computations. However, the alternating optimization framework of DMRG is prone to convergence to local minima for certain types of systems [37]. To address this issue, several variations of the DMRG method have been proposed, such as subspace enrichment techniques that increase the rank during optimization [7]. Despite these enhancements, the DMRG algorithm and its variants remain unreliable for certain systems. Furthermore, the original DMRG formulation was designed to compute only the lowest eigenpair of $A$. Extensions for computing multiple eigenpairs often involve sequential deflation, i.e., only one eigenvector is computed at a time, or block low-rank representations [6, 18, 14, 15], which may require higher ranks than are necessary to represent the eigenvectors individually, leading to inefficiencies.
@@
In this paper, we examine the application of the Lanczos method and polynomial-filtered subspace iteration, two classical iterative methods, for computing several extreme eigenpairs of $A$. Both methods employ a low-rank tensor representation of $A$, such as the CP representation in (2) or TT-matrix representation, and represent all vectors in a compatible low-rank tensor format. A key challenge in applying these iterative methods is that standard operations, such as matrix-vector products and vector addition, can increase tensor rank, necessitating truncation that introduces potentially large errors. Low-rank variants of Krylov methods have been studied in [5, 19, 20, 26]. The Lanczos method constructs an approximate orthonormal basis of a Krylov subspace, which we show are often difficult to represent accurately using low-rank tensor formats, even when the eigenvectors themselves have low-rank structure. We further show that the truncation errors introduced while constructing the approximate Krylov subspace result in a significant loss of accuracy in the Lanczos method. In contrast, polynomial-filtered subspace iteration directly uses approximate eigenvectors (Ritz vectors) as a subspace basis, avoiding the need for generating an orthonormal basis for a Krylov subspace. Through analysis and numerical experiments, we demonstrate that subspace iteration is much more robust to rank-truncation errors than the Lanczos method. To the best of our knowledge, the present paper is the first study of polynomial-filtered subspace iteration with low-rank tensor compression. Approximate power iteration using sparse matrix and vector representations has been studied in [21, 38].
@@
The paper is organized as follows. In Section 2, we briefly review the tensor train (TT) representation for high-dimensional tensors, including basic operations and their computational costs. Section 3 describes the Lanczos algorithm, originally developed in [5], for solving low-rank tensor eigenvalue problems. In Section 4, we introduce a low-rank variant of Chebyshev filtered subspace iteration (ChebFSI). Section 5 discusses the convergence properties of these low-rank subspace projection methods. Finally, in Section 7, we compare the convergence of the Lanczos and ChebFSI methods with varying amounts of rank truncation, and provide numerical demonstrations our theoretical convergence results.
@@ 2 Low-rank tensor representation
The vector space $\mathbb{C}^{n_{1}\cdots n_{d}}$, on which (2) acts, is naturally isomorphic to the tensor space $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$, as tensors can be placed in one-to-one correspondence with vectors by reshaping. We refer to $d$ as the dimension and $n_{k}$ ($k=1,2,\ldots,d$) as the mode sizes of the tensor. The total number of entries in the tensor grows exponentially with the dimension, making the storage and computation of eigenvectors for (2) infeasible on modern computing devices. To reduce the number of degrees of freedom, we seek eigenvectors represented in a low-rank tensor format. A low-rank tensor format factorizes a tensor $v$ into a network of low-dimensional tensors, enabling more efficient storage and computation. Several tensor formats are available, including tensor train (TT) [24], hierarchical Tucker [8], canonical polyadic (CP), and Tucker [17]. In this paper, we focus on tensors and operators represented in the TT format, also known as the matrix product state (MPS) in the physics literature [23]. The algorithms discussed in the subsequent sections can be readily adapted to other tensor formats. In the remainder of this section, we provide a brief overview of the TT format and its associated arithmetic operations.
@@ 2.1 Tensor train format
(3) $v(i_{1},i_{2},\ldots,i_{d})=C_{1}(i_{1})C_{2}(i_{2})\cdots C_{d}(i_{d}),$
where each $C_{k}$ is a $r_{k-1}\times n_{k}\times r_{k}$ tensor, referred to as a TT-core. Here, $C_{k}(i_{k})$ is interpreted as a $r_{k-1}\times r_{k}$ matrix with $r_{0}=r_{d}=1$. Thus each element of $v$ is represented as a product of $d$ matrices, hence the name matrix product state (MPS) given to this representation in physics literature. If we assume the $r_{k}$ are all equal to $r$ and the mode size $n_{k}$ are all equal to $n$, the decomposition (3) has $\mathcal{O}(dnr^{2})$ degrees of freedom. Importantly, this number scales linearly with the number of dimensions $d$, provided the parameter $r$ remains constant. The vector $\bm{r}=(1,r_{1},\ldots,r_{d-1},1)$ containing the minimal $r_{k}$ needed to represent $v$ is called the TT-rank of $v$ and the efficiency of the TT representation (3) relies on keeping the TT-rank as small as possible. To apply the Lanczos method or subspace iteration for computing eigenvectors of (2) in the TT format (3), we first review the essential arithmetic operations on TTs.
@@ 2.1.1 Addition of tensors in TT format
Given TTs $v(i_{1},\ldots,i_{d})=C_{1}(i_{1})\cdots C_{d}(i_{d})$ and $w(i_{1},\ldots,i_{d})=D_{1}(i_{1})\cdots D_{d}(i_{d})$ with compatible mode sizes, the sum $v+w$ can be represented in the TT format by concatenating the cores of $v$ and $w$. The dimensions of the concatenated cores equal to the sum of the dimensions of $C_{k}$ and $D_{k}$. To control the TT-ranks, sums are typically followed by a rank-truncation procedure or directly approximated with a low-rank tensor, as described in Section 2.1.3.
@@ 2.1.2 TT matrices and matrix-vector products
The matrix (2) is expressed in a low-rank format known as the canonical polyadic (CP) decomposition. This Kronecker product structure enables the efficient application of each summand to a TT. By subsequently approximating the sum with a low-rank TT as described in Section 2.1.3, we obtain an efficient method for computing an approximate matrix-vector product between (2) and a TT. Alternatively, any operator expressed in the CP format (2) can also be written as a TT-matrix [18, Section 2.3], also referred to as matrix product operator (MPO) in the physics literature, as
(4) $A(i_{1}\cdots i_{d},j_{1}\cdots j_{d})=A_{1}(i_{1},j_{1})A_{2}(i_{2},j_{2})\cdots A_{d}(i_{d},j_{d}),$
@@ 2.1.3 TT rank-truncation
Common operations, such as addition and matrix-vector products between TTs, as described above, result in TTs with increased rank. To maintain low-rank vectors and hence computationally efficient algorithms, the result of these operations are typically approximated with a TT of smaller rank. A straightforward approach to computing such approximations involves first constructing the TT with larger ranks and then recursively applying SVD-based truncation with the TT-SVD algorithm [24].
We denote by $\mathfrak{T}_{\bm{r}}(w)$ the result of truncating $w$ to rank-$\bm{r}$ with TT-SVD. In each step, if the rank $r_{j}$ is determined by discarding all singular values smaller than $\delta=\epsilon/\sqrt{d-1}$ for some $\epsilon>0$, then the error introduced by truncating $w$ to rank-$\bm{r}$ satisfies
(5) $\|\mathfrak{T}_{\bm{r}}(w)-w\|_{F}\leq\epsilon.$
The computational cost of applying $\mathfrak{T}_{\bm{r}}$ to a TT with cores of dimension $\bm{s}$ is $\mathcal{O}(dns^{3})$, assuming $\bm{s}$ is constantly equal to $s$ and the mode sizes are constantly equal to $n$. If the cores of $w$ are obtained from matrix multiplication between a TT-matrix $A$ with ranks equal to $r_{A}$ and a TT $v$ with ranks equal to $r_{v}$, then the cost of TT-SVD truncation for $w$ is $\mathcal{O}(dnr_{A}^{3}r_{v}^{3})$ which can be a computational bottleneck in the iterative methods described in subsequent sections if either $r_{A}$ or $r_{v}$ are moderately large. Additionally, when computing linear combinations of several TTs, which is required in the iterative methods described below, TT-SVD truncation is often inserted after each addition in order to control the rank of intermediate TTs. This process can become costly and inaccurate as the number of summands increases, since many truncations are required and truncation errors can accumulate. Moreover, the order in which the summands are added can significantly affect the final result.
@@
Other approaches to approximating sums and matrix-vector products with low-rank TTs avoid constructing intermediate TTs with large ranks, offering greater computational efficiency at the expense of some accuracy. Examples include randomized algorithms [1], projections onto a suitable TT tangent space [33, Section 4.3], and variational methods. Unlike recursively applying the TT-SVD truncation, these approaches are not influenced by the ordering of the summands when approximating linear combinations. We demonstrate in Section 7 that the convergence of the polynomial filtered subspace iteration is robust to truncation errors. This robustness makes it advantageous to employ less accurate but faster TT truncation procedures, such as randomized methods, tangent space projection, variational methods, or some combination thereof, which can significantly accelerate computations without compromising the overall convergence of the algorithm.
2.1.4 Orthogonal projection onto TT tangent space
The collection of all TTs with rank-$\bm{r}$
(6) $\mathcal{M}_{\bm{r}}=\{v\in\mathbb{C}^{n_{1}\times\cdots\times n_{d}}\mid\text{TT-rank}(v)=\bm{r}\},$
forms a smooth submanifold of $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$ [13]. Associated with any $v\in\mathcal{M}_{\bm{r}}$ is a tangent space denoted by $T_{v}\mathcal{M}_{\bm{r}}$, which is a vector subspace of $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$. Such tangent spaces, typically used for dynamical low-rank approximation of tensor differential equations [22, 11], are a convenient set of vector subspaces in which we can approximate the linear algebra operations arising in the inexact subspace projection methods described below. Given the TT representation (3) of $v$, any element of the tangent space can be written (non-uniquely) as
(7) $\delta v=\delta C_{1}C_{2}\cdots C_{d}+C_{1}\delta C_{2}C_{3}\cdots C_{d}+\cdots+C_{1}\cdots C_{d-1}\delta C_{d},$
2.1.5 Inner product of tensors in TT format
3 Lanczos method
(8) $\mathcal{K}(A,v_{0};m)=\text{span}\{v_{0},Av_{0},...,A^{m-1}v_{0}\}.$
(9) $AV_{m}=V_{m}T_{m}+fe_{m}^{\text{T}},$
(10) $\beta_{j+1}v_{j+1}=Av_{j}-\alpha_{j}v_{j}-\beta_{j}v_{j-1},$
3.1 Lanczos with rank truncation
(11) $\tilde{T}_{m}=R^{-\text{T}}\left(V_{m}AV_{m}\right)R^{-1},$
Algorithm 1 Lanczos with rank truncation
1:$A\in\mathbb{C}^{n_{1}\cdots n_{d}\times n_{1}\cdots n_{d}}$, $v_{0}\in\mathbb{C}^{n_{1}\cdots n_{d}}$ initial low-rank tensor with norm $1$, $m$ number of Lanczos iterations
2:$(v_{0},v_{1},\ldots,v_{m-1})$ Approximate Krylov basis vectors.
3:$v_{1}=\mathfrak{T}_{\bm{r}}\left(Av_{0}\right)$
4:$\alpha_{0}=\left\langle v_{1},v_{0}\right\rangle$
5:$v_{1}=\mathfrak{T}_{\bm{r}}\left(v_{1}-\alpha_{0}v_{0}\right)$
6:for $j=1$ to $m-1$ do
7:     $\beta_{j}=\|v_{j}\|$
8:     $v_{j}=v_{j}/\beta_{j}$
9:     $v_{j+1}=\mathfrak{T}_{\bm{r}}\left(Av_{j}\right)$
10:     $\alpha_{j}=\left\langle v_{j+1},v_{j}\right\rangle$
11:     $v_{j+1}=\mathfrak{T}_{\bm{r}}\left(v_{j+1}-\alpha_{j}v_{j}-\beta_{j}v_{j-1}\right)$
12:end for

### v3

# Inexact subspace projection methods for low-rank tensor eigenvalue problems

@@ Abstract
We propose inexact subspace iteration for solving high-dimensional eigenvalue problems with low-rank structure. Inexactness stems from low-rank compression, enabling efficient representation of high-dimensional vectors in a low-rank tensor format. A primary challenge in these methods is that standard operations, such as matrix-vector products and linear combinations, increase tensor rank, necessitating rank truncation and hence approximation. We compare the proposed methods with an existing inexact Lanczos method with low-rank compression. This method constructs an approximate orthonormal Krylov basis, which is often difficult to represent accurately in low-rank tensor formats, even when the eigenvectors themselves exhibit low-rank structure. In contrast, inexact subspace iteration uses approximate eigenvectors (Ritz vectors) directly as a subspace basis, bypassing the need for an orthonormal Krylov basis. Our analysis and numerical experiments demonstrate that inexact subspace iteration is much more robust to rank-truncation errors compared to the inexact Lanczos method. We also demonstrate that rank-truncated subspace iteration can converge for problems where the DMRG method stagnates. Furthermore, the proposed subspace iteration methods do not require a Hermitian matrix, in contrast to Lanczos and DMRG, which are designed specifically for Hermitian matrices.
journal: Arxiv
@@ 1 Introduction
of a matrix $A\in\mathbb{C}^{(n_{1}\cdots n_{d})\times(n_{1}\cdots n_{d})}$ that can be written as a sum of Kronecker products
$A=\sum_{j=1}^{r_{A}}A_{1}^{(j)}\otimes A_{2}^{(j)}\otimes\cdots\otimes A_{d}^{(j)},$ (2)
@@
where $r_{A}$ is the separation (or canonical polyadic) rank of $A$ and $A_{i}^{(j)}\in\mathbb{C}^{n_{i}\times n_{i}}$ for all $i=1,2,\ldots,d$ and $j=1,\ldots,r_{A}$. Examples of (2) appear naturally in quantum many-body physics and arise in many other applications from the discretization of partial differential operators on separable domains. The number of entries in $A$ and its eigenvectors $\psi\in\mathbb{C}^{n_{1}\cdots n_{d}}$ grows exponentially fast with the parameter $d$, making classical numerical methods too expensive when directly applied to $A$, even for moderate $d$. Given the tensor product structure of the operator (2), it is natural to seek eigenvectors in a compact form commensurate with the tensor product structure of the operator $A$. Many tensor networks (or formats) are available for representing operators and vectors [9], offering significant reductions in the degrees of freedom required, provided the objects can be approximated with a network of relatively small tensor rank. For quantum many-body Hamiltonians with separated eigenvalues and local interactions on 1D geometries, it is well-known that the eigenvectors corresponding to the smallest eigenvalues admit accurate low-rank tensor train (or matrix product state) approximations, thanks to the area law [12]. For other applications, e.g., the spatial discretization of high-dimensional elliptic operators, it has been demonstrated that eigenvectors can be well-approximated in low-rank hierarchical tensor formats [3, 10].
@@
A widely used approach for computing the extreme eigenpairs of a Hermitian matrix $A$, expressed in the form (2), is the density matrix renormalization group (DMRG) method [30, 31, 36], originally developed in the physics community. The DMRG method represents the approximate eigenvector as a tensor network and optimizes one (or two) low-dimensional tensors in the network at a time by projecting $A$ onto subspaces defined by the remaining tensors. This optimization process cycles through all tensors in the network, often multiple times, until convergence. The method achieves computational efficiency by reducing the high-dimensional eigenvalue problem to a sequence of low-dimensional eigenvalue computations. However, the alternating optimization framework of DMRG is prone to convergence to local minima for certain types of systems [37]. To address this issue, several variations of the DMRG method have been proposed, such as subspace enrichment techniques that increase the rank during optimization [7]. Despite these enhancements, the DMRG algorithm and its variants remain unreliable for certain systems. Furthermore, the original DMRG formulation was designed to compute only the lowest eigenpair of $A$. Extensions for computing multiple eigenpairs often rely on sequential deflation, where eigenvectors are computed one at a time, or on block low-rank representations [6, 18, 14, 15]. In our experience, the latter can require higher ranks than are needed to represent the eigenvectors individually, leading to inefficiencies. Also, DMRG is based on an energy minimization principle that requires $A$ to be Hermitian. This requirement is not imposed by the methods proposed in this paper.
@@
We propose an inexact polynomial-filtered subspace iteration for computing several extreme eigenpairs of $A$. We employ a low-rank tensor representation of $A$, such as the CP representation in (2) or TT-matrix representation, and represent all vectors in a compatible low-rank tensor format. A key challenge in applying this iterative method is that standard operations, such as matrix-vector products and vector addition, often increase tensor rank, necessitating truncation that introduces potentially large errors. We provide direct comparisons of the proposed subspace iteration with low-rank variants of Krylov methods for symmetric matrices, i.e., Lanczos method, proposed in [5, 19, 20, 26]. These methods construct an approximate orthonormal basis of a Krylov subspace, which we show are often difficult to represent accurately using low-rank tensor formats, even when the eigenvectors themselves have low-rank structure. We further show that the truncation errors introduced while constructing the approximate Krylov subspace result in a significant loss of accuracy in the Krylov method. In contrast, polynomial-filtered subspace iteration directly uses approximate eigenvectors (Ritz vectors) as a subspace basis, avoiding the need for generating an orthonormal basis for a Krylov subspace. Through analysis and numerical experiments, we demonstrate that subspace iteration is much more robust to rank-truncation errors than the Lanczos method. To the best of our knowledge, the present paper is the first study of polynomial-filtered subspace iteration with low-rank tensor compression. Approximate power iteration using sparse matrix and vector representations has been studied in [21, 38].
@@
The paper is organized as follows. In section 2, we briefly review the tensor train (TT) representation for high-dimensional tensors and basic operations involving TTs. In section 3, we introduce tensor train Chebyshev filtered subspace iteration with rank-truncation. In section 4 we provide a convergence analysis that gives an upper bound on the amount of acceptable truncation error in TT subspace iteration. In section 5 we discuss the computational cost of TT subspace iteration and different algorithms for performing rank-truncation. Finally, in section 6, we provide numerical comparisons between the proposed subspace iteration methods and Lanczos and DMRG, provide numerical demonstrations of our theoretical convergence results, and compare rank-truncation algorithms.
@@ 2 Low-rank tensor representation
The vector space $\mathbb{C}^{n_{1}\cdots n_{d}}$, on which (2) acts, is naturally isomorphic to the tensor space $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$, as tensors can be placed in one-to-one correspondence with vectors by reshaping. We refer to $d$ as the dimension and $n_{k}$ ($k=1,2,\ldots,d$) as the mode sizes of the tensor. The total number of entries in the tensor grows exponentially with the dimension, making the storage and computation of eigenvectors for (2) infeasible on modern computing devices. To reduce the number of degrees of freedom, we seek eigenvectors represented in a low-rank tensor format. A low-rank tensor format factorizes a tensor $v$ into a network of low-dimensional tensors, enabling more efficient storage and computation. Several tensor formats are available, including tensor train (TT) [24], hierarchical Tucker [8], canonical polyadic (CP), and Tucker [17]. In this paper, we focus on tensors and operators represented in the TT format, also known as the matrix product state (MPS) in the physics literature [23]. The algorithms discussed in the subsequent sections can be readily adapted to other tensor formats. In the remainder of this section, we provide a brief overview of the TT format and arithmetic operations involving TTs.
@@ 2.1 Tensor train format
$v(i_{1},i_{2},\ldots,i_{d})=C_{1}(i_{1})C_{2}(i_{2})\cdots C_{d}(i_{d}),$ (3)
where each $C_{k}$ is a $r_{k-1}\times n_{k}\times r_{k}$ tensor, referred to as a TT-core. Here, $C_{k}(i_{k})$ is interpreted as a $r_{k-1}\times r_{k}$ matrix with $r_{0}=r_{d}=1$. Thus each element of $v$ is represented as a product of $d$ matrices, hence the name matrix product state (MPS) given to this representation in physics literature. If we assume the $r_{k}$ are all equal to $r$ and the mode size $n_{k}$ are all equal to $n$, the decomposition (3) has $\mathcal{O}(dnr^{2})$ degrees of freedom. Importantly, this number scales linearly with the number of dimensions $d$, provided the parameter $r$ remains constant. The vector $\bm{r}=(1,r_{1},\ldots,r_{d-1},1)$ containing the minimal $r_{k}$ needed to represent $v$ is called the TT-rank of $v$ and the efficiency of the TT representation (3) relies on keeping the TT-rank as small as possible. To apply the Lanczos method or subspace iteration for computing eigenvectors of (2) in the TT format (3), we first review the essential arithmetic operations on TTs. A key challenge of iterative methods with TTs is that standard linear algebra operations, such as matrix-vector products and linear combinations, increase TT rank and hence are often followed by rank-truncation to control computational cost. Such rank-truncation is the source of inexactness in the TT algorithms to follow.
2.1.1 Inner product of tensors in TT format
@@ 2.1.2 Addition of tensors in TT format
Given TTs $v(i_{1},\ldots,i_{d})=C_{1}(i_{1})\cdots C_{d}(i_{d})$ and $w(i_{1},\ldots,i_{d})=D_{1}(i_{1})\cdots D_{d}(i_{d})$ with compatible mode sizes, the sum $v+w$ can be represented in the TT format by concatenating the cores of $v$ and $w$. The dimensions of the concatenated cores, and hence the resulting TT rank, is equal to the sum of the dimensions of $C_{k}$ and $D_{k}$. To control the TT-ranks, sums are typically followed by a rank-truncation procedure or directly approximated with a low-rank tensor, as described in section 2.1.4.
@@ 2.1.3 TT matrices and matrix-vector products
The matrix (2) is represented in a low-rank format known as the canonical polyadic (CP) decomposition. This Kronecker product structure allows each summand to be applied efficiently to a TT. After applying all summands, the resulting sum can be approximated by a low-rank TT using the truncation procedure described in section 2.1.4. This yields an efficient method for computing approximate matrix–vector products between (2) and a TT.
Alternatively, any operator expressed in the CP format (2) can be written as a TT-matrix [18, Section 2.3], also referred to as matrix product operator (MPO) in the physics literature, as
$A(i_{1}\cdots i_{d},j_{1}\cdots j_{d})=A_{1}(i_{1},j_{1})A_{2}(i_{2},j_{2})\cdots A_{d}(i_{d},j_{d}),$ (4)
@@ 2.1.4 TT rank-truncation
Common operations, such as addition and matrix-vector products between TTs, as described above, result in TTs with increased rank. To maintain low-rank vectors and hence computationally efficient algorithms, the result of these operations are typically approximated with a TT of smaller rank. A straightforward approach to computing rank $\bm{r}$ TT approximations of sums and matrix-vector products involves first constructing a TT with larger ranks and then recursively applying SVD-based truncation with the TT-SVD algorithm [24]. The user can select a maximum allowable rank $\bm{r}$ or a desired accuracy $\epsilon>0$. In the latter case the TT-SVD algorithm selects the rank $r_{j}$ by discarding all singular values smaller than $\delta=\epsilon/\sqrt{d-1}$ of a relevant matrix, so that the error introduced by truncating $w$ to rank-$\bm{r}$ satisfies
$\|\mathfrak{T}_{\bm{r}}(w)-w\|_{F}\leq\epsilon.$ (5)
@@
Other approaches to approximating sums and matrix-vector products with low-rank TTs avoid constructing intermediate TTs with large ranks, offering greater computational efficiency at the expense of some accuracy. Examples include randomized algorithms [1], projections onto a suitable TT tangent space [33, Section 4.3], and variational methods. Such approaches are discussed further in section 5.1 and section 5.2. In the development of the following algorithms, we use $\mathfrak{T}_{\bm{r}}$ to denote a rank-$\bm{r}$ truncation operator without specifying a particular truncation algorithm. Computational costs of different truncation algorithms are discussed in section 5 and comparisons are presented in our numerical experiments in section 6.5.